OpenStudy (anonymous):

A roast turkey is taken from an oven when its temperature has reached 185°F and is placed on a table in a room where the temperature is 75°F. (Round your answers to the nearest whole number.)

OpenStudy (anonymous):

@kimmy2727 so whats the actual question?

OpenStudy (anonymous):

(a) If the temperature of the turkey is 150°F after half an hour, what is the temperature after 50 minutes? T(50) =

OpenStudy (anonymous):

and

OpenStudy (anonymous):

(b) When will the turkey have cooled to 105°? t = min

OpenStudy (anonymous):

for (a)..Find the equation relating time and temp: intercept = 185 slope + (185-150)/(0-30) = -35/30 = -7/6 then...Equation: f(t) = (-7/6)t+185 what is its temperature after 50 min? f(50) = (-7/6)(50)+185 = ?

OpenStudy (anonymous):

& for (b)...make 105 the subject of f(t) = (-7/6)t + 185 eg. 105 = (-7/6)t+185 (-7/6)t = -80 t = ? minutes.